# ESA: Annotation-Efficient Active Learning for Semantic Segmentation

### v1

@@ 3.5 Algorithm
1 Require: Source dataset with annotation $(\mathbf{I}_{s},\mathbf{Y}_{s})$, target dataset without label $\mathbf{I}_{t}$, total iteration $N$, selection rounds $S$, per-round budget $b$, and hyperparameters: $\tau$, $\alpha$.
@@ 4 Experiments
Implementation details. We adapt DeepLab-v3+ and ResNet-101 as model backbones on Tesla V100 GPU using PyTorch. Our models are optimized using the SGD optimizer with a momentum of 0.9, weight decay of 0.0005, and poly learning rate schedule with the initial learning rate set to 0.00025. The training process is stopped after 40k iterations with batch size 2. Both the source and target datasets are resized to a uniform dimension of, 512$\times$512 to maintain consistency during training. For Superpixel-based Annotation (SA), we set the parameters k to 1000 and compactness to 0.1. Furthermore, we have reimplemented RIPU [31] as our baseline.
@@ 4.1 Main Results
Table 1: Comparison with previous results on task COCO $\rightarrow$ VOC. The detailed information is presented in Sec  1. ‘*’ means we re-implement previous method for COCO $\rightarrow$ VOC dataset.
@@
Source Only 91.16 86.18 37.25 80.09 57.99 45.64 87.43 78.63 86.29 28.94 70.66 37.46 84.28 64.61 71.83 83.36 38.25 76.19 40.19 72.00 71.91 66.21
WDA(Image) [18] 91.26 82.7 30.73 80.13 66.12 49.93 92.77 69.17 84.99 33.21 88.82 52.66 80.42 81.89 75.79 81.31 41.60 79.53 50.29 82.84 63.87 69.53
WDA(Point) [18] 91.76 81.03 32.31 76.71 67.24 53.29 93.07 72.23 87.38 33.11 87.71 58.32 79.95 79.3 73.98 83.3 44.31 81.13 53.61 81.38 61.6 70.13
RIPU (PA 40) [31] 93.60 89.74 51.65 83.48 71.81 65.96 91.32 83.77 90.46 28.09 72.77 57.70 82.14 75.38 82.16 84.94 35.86 73.56 44.99 78.97 72.07 71.93
Ours (SA 40) 92.61 86.82 42.85 83.37 73.87 59.96 89.89 84.07 87.85 36.15 83.11 59.71 81.12 77.86 80.62 83.93 54.40 82.02 47.66 82.19 74.31 73.54
RIPU (PA 1000) [31] 93.67 87.94 55.08 83.53 68.32 61.67 90.37 81.86 87.64 29.85 74.09 55.76 82.87 75.15 78.64 85.35 52.60 68.02 51.01 82.15 75.17 72.42
Ours (SA 1000) 93.87 88.73 54.14 87.01 74.84 72.20 88.05 85.43 90.49 35.33 73.07 54.53 86.53 72.94 81.10 85.85 50.42 82.99 51.74 80.19 75.52 74.52
Ours (EA 33) 93.11 83.59 42.44 88.84 67.34 67.90 91.16 82.33 89.94 28.93 77.98 56.45 84.50 77.15 79.97 85.10 52.95 79.12 47.68 78.85 75.77 72.91
RIPU (PA, 5767)  [31] 93.13 83.18 46.9 84.62 69.47 61.77 87.68 84.8 88.52 36.19 80.17 57.86 84.22 80.54 73.08 84.82 45.70 82.36 46.17 73.01 76.53 72.41
RIPU (RA, 640)  [31] 93.96 91.80 56.53 87.25 74.78 68.43 84.80 85.64 88.69 37.25 82.05 48.85 84.90 78.14 81.96 86.03 53.19 81.90 49.04 78.47 74.76 74.69
Ours (ESA, 102) 93.23 88.49 42.10 89.22 75.10 70.01 88.57 84.49 88.75 36.99 72.57 56.25 84.24 77.92 82.65 84.16 53.91 79.67 49.05 83.89 75.27 74.12
Fully Supervised 93.63 89.49 59.32 82.60 72.89 61.93 88.23 84.77 89.18 30.72 81.13 57.26 85.10 76.90 81.84 85.31 51.07 75.48 46.41 78.01 71.72 73.48
@@ 4.3 Ablation Study
Table 3: Experiments on different active selection methods. Comparison of active selection methods: random selection (RAND), entropy-based selection (ENT), softmax confidence based selection (SCONF), and our method as described in Sec. 3.3.

### v2

@@ 3.5 Algorithm
1 Require: Source dataset with annotation $(\mathbf{I}_{s},\mathbf{Y}_{s})$, target dataset without label $\mathbf{I}_{t}$, total iteration NN, selection rounds $S$, per-round budget $b$, and hyperparameters: $\tau$, $\alpha$.
@@ 4 Experiments
Implementation details. We adapt DeepLab-v3+ and ResNet-101 as model backbones on Tesla V100 GPU using PyTorch. Our models are optimized using the SGD optimizer with a momentum of 0.9, weight decay of 0.0005, and poly learning rate schedule with the initial learning rate set to 0.00025. The training process is stopped after 40k iterations with batch size 2. Both the source and target datasets are resized to a uniform dimension of, $\times512$ to maintain consistency during training. For Superpixel-based Annotation (SA), we set the parameters k to 1000 and compactness to 0.1. Furthermore, we have reimplemented RIPU [31] as our baseline.
@@ 4.1 Main Results
Table 1: Comparison with previous results on task COCO $\rightarrow$ VOC. Detailed results are presented in Sec 1. ‘*’ denotes re-implemented baselines for COCO $\rightarrow$ VOC task. Abbreviated class names: back. = background, aero. = aeroplane, bott. = bottle, moto. = motorbike, pers. = person, plant = potted plant, tv = TV/monitor. Full class names: bicycle, bird, boat, bus, car, cat, chair, cow, dining table, dog, horse, sheep, sofa, train.
@@
Source Only 91.16 86.18 37.25 80.09 57.99 45.64 87.43 78.63 86.29 28.94 70.66
37.46 84.28 64.61 71.83 83.36 38.25 76.19 40.19 72.00 71.91 66.21
WDA(Image) [18] 91.26 82.7 30.73 80.13 66.12 49.93 92.77 69.17 84.99 33.21 88.82
52.66 80.42 81.89 75.79 81.31 41.60 79.53 50.29 82.84 63.87 69.53
WDA(Point) [18] 91.76 81.03 32.31 76.71 67.24 53.29 93.07 72.23 87.38 33.11 87.71
58.32 79.95 79.3 73.98 83.3 44.31 81.13 53.61 81.38 61.6 70.13
RIPU (PA 40) [31] 93.60 89.74 51.65 83.48 71.81 65.96 91.32 83.77 90.46 28.09 72.77
57.70 82.14 75.38 82.16 84.94 35.86 73.56 44.99 78.97 72.07 71.93
Ours (SA 40) 92.61 86.82 42.85 83.37 73.87 59.96 89.89 84.07 87.85 36.15 83.11
59.71 81.12 77.86 80.62 83.93 54.40 82.02 47.66 82.19 74.31 73.54
RIPU (PA 1000) [31] 93.67 87.94 55.08 83.53 68.32 61.67 90.37 81.86 87.64 29.85 74.09
55.76 82.87 75.15 78.64 85.35 52.60 68.02 51.01 82.15 75.17 72.42
Ours (SA 1000) 93.87 88.73 54.14 87.01 74.84 72.20 88.05 85.43 90.49 35.33 73.07
54.53 86.53 72.94 81.10 85.85 50.42 82.99 51.74 80.19 75.52 74.52
Ours (EA 33) 93.11 83.59 42.44 88.84 67.34 67.90 91.16 82.33 89.94 28.93 77.98
56.45 84.50 77.15 79.97 85.10 52.95 79.12 47.68 78.85 75.77 72.91
RIPU (PA, 5767)  [31] 93.13 83.18 46.9 84.62 69.47 61.77 87.68 84.8 88.52 36.19 80.17
57.86 84.22 80.54 73.08 84.82 45.70 82.36 46.17 73.01 76.53 72.41
RIPU (RA, 640)  [31] 93.96 91.80 56.53 87.25 74.78 68.43 84.80 85.64 88.69 37.25 82.05
48.85 84.90 78.14 81.96 86.03 53.19 81.90 49.04 78.47 74.76 74.69
Ours (ESA, 102) 93.23 88.49 42.10 89.22 75.10 70.01 88.57 84.49 88.75 36.99 72.57
56.25 84.24 77.92 82.65 84.16 53.91 79.67 49.05 83.89 75.27 74.12
Fully Supervised 93.63 89.49 59.32 82.60 72.89 61.93 88.23 84.77 89.18 30.72 81.13
57.26 85.10 76.90 81.84 85.31 51.07 75.48 46.41 78.01 71.72 73.48
@@ 4.3 Ablation Study
Table 3: Experiments on different active selection methods. Comparison of active selection methods: random selection (RAND), entropy-based selection (ENT), softmax confidence based selection (SCONF), and our method as described in Sec. $3.$